# A Generalized Theory for Optical Cooling of a Trapped Atom with Spin

### v2

@@ Abstract
Cooling atoms to the ground-state of optical tweezers is becoming increasingly important for high-fidelity imaging, cooling, and molecular assembly. While extensive theoretical work has been conducted on cooling in free space, fewer studies have focused on cooling in bound states. In this work, we present a unified formalism for optical cooling mechanisms in neutral atom tweezers, including resolved and unresolved sideband cooling with different trapping potentials, polarization gradient cooling, gray molasses cooling, $\Lambda$-enhanced gray molasses cooling, and Raman sideband cooling. We perform simulations and demonstrate good agreement with a simplified spin model. We derive and discuss the fundamental limits of each cooling mechanism and propose new strategies for achieving ground-state cooling in optical tweezers. Our findings provide valuable insights into optimizing cooling schemes for neutral atoms in optical tweezers, paving the way for minimizing thermal decoherence in Rydberg and molecular gates and improving efficiencies of molecular assembly.
@@ I Introduction
Laser cooling of bound neutral atoms has recently gained renewed significance in experiments involving tweezer arrays of atoms and molecules [1]. In systems of Rydberg atoms [2] or molecules [3, 4, 5] with dipolar coupling, thermal motion limits the coherence of interactions. Molecular assembly [6, 7] and few-body physics [8, 9] studies require high-fidelity ground-state preparation of multiple atoms.
@@
A wide range of optical cooling techniques have been developed for trapped atoms, illustrated in Fig. 1. Single-photon cooling schemes, such as sideband cooling, have been used for atoms with narrow lines for high-fidelity imaging and ground-state preparation [10, 11, 12]. Two-photon schemes, including polarization gradient (PG) [13, 14, 15, 16], gray molasses (GM) and $\Lambda$-enhanced gray molasses ($\Lambda$-GM) [17, 18, 19, 20, 21], and Raman-sideband cooling (RSB) [22, 23, 24, 25, 26], have been successfully applied for imaging and cooling of atoms without narrow lines. While extensive theoretical work has been developed for sideband cooling [27, 28, 29], many of the other cooling schemes for bound atoms are still interpreted through their free-space pictures, for example in Sisyphus cooling where an atom moves through a changing polarization [13]. However, this picture breaks down for a tightly trapped atom in the Lamb-Dicke regime, where the atom wavefunction is smaller than the wavelength.
@@
In this work, we develop a generalized optical cooling model for bound atoms with spin in a single or counter-propagating beam configuration with different polarizations. We validate these models using master equation simulations that include the ground and excited spins and the harmonic oscillator states. Our analysis reveals a consistent picture across various cooling schemes, shown in Fig. 1. Cooling occurs when $|n\rangle$ and $|n-1\rangle$ are Raman coupled for different spin states and brought into resonance using a combination of light shifts (PG, GM, $\Lambda$-GM) and two-photon detunings (RSB).
@@
Sideband cooling has been extensively studied in previous works [30, 31], but cooling of a bound atom with spin has been much more limited. Previous work on cooling bound atoms with spin has primarily focused on specific cases, such as $J\!\!=\!\!1/2$ to $J^{\prime}\!\!=\!\!3/2$ transitions in counter-propagating orthogonal linear light [32] or linear light with standing waves [33]. We develop a new formalism for arbitrary spin and polarization, with a particular emphasis on the experimentally relevant $\sigma_{+}-\sigma_{-}$ configuration. Notably, while these previous theoretical studies found a fundamental limit of $\langle n\rangle\approx 1$, we demonstrate that schemes like GM and $\Lambda$-GM can be used for high ground-state preparation when considering different polarization and spin configurations.
@@
Figure 1: (a) Laser cooling of a trapped atom with ground and excited state spin $F$ and $F^{\prime}$ in a changing polarization from counter-propagating beams. (b) Single photon cooling schemes such as resolved and unresolved sideband cooling drive population from the ground-state to the excited state with one reduced motional quanta $n-1$. In the Lamb-Dicke regime, the recoil heating is smaller than the trapping frequency and the population mostly returns to $n-1$, resulting in net cooling. (c) Two-photon cooling schemes like polarization gradient (PG), gray molasses (GM), and $\Lambda$-enhanced gray molasses ($\Lambda$-GM) drive $n\rightarrow n-1$ between two ground-state spins with a two-photon Raman transition. (d) Temperature scale for cooling schemes. A wide variety of cooling schemes have been used to trap, cool, and image tightly trapped atoms.
@@
A recent experimental result achieved a record-high imaging fidelity of 99.96% for lithium using $\Lambda$-enhanced GM cooling with significant ground-state preparation, despite lithium having the largest recoil heating [21]. This raises the question of whether spin cooling techniques can be further optimized for ground-state preparation. The unified model also allows us to explore novel cooling schemes that transcend the conventions of any single conventional cooling technique.
@@
The paper is structured as follows. Section II develops the general master equation for spin cooling and derive effective ground-state spin operators by adiabatically eliminating the excited state. Section III applies this formalism to sideband cooling, including cases with different excited state trapping frequencies. Section IV develops a formalism for spin cooling and applies it to PG, GM, $\Lambda$-GM, and RSB cooling. Each section compares exact simulations to simplified models and contrasts various cooling techniques, with the goal of creating a unified picture that will lead to more coherent interactions and higher fidelity ground-state preparation.
@@ II.1 Master equation for laser cooling
In this section, we develop a master equation for an atom confined in a tight harmonic trap and interacting with a semi-classical electric field. In a semi-classical approximation, we take the expectation of the electric field but keep the atom position $\hat{x}$ as a quantum operator [34, 35]. The electric field for a single laser frequency $\omega_{L}$ is
$\bm{E}(\hat{x},t)=E_{0}\,{\bm{\mathcal{\epsilon}}}(\hat{x})\,e^{-i\omega_{L}t}+h.c..$ (1)
$E_{0}$ is a real field amplitude and the last term is the hermitian conjugate. The complex polarization and phase are contained within $\bm{\epsilon}(\hat{x})$ and can describe a traveling wave or a standing wave with various polarization configurations.
@@
$\hat{H}=\nu\hat{a}^{\dagger}\hat{a}+\left(\Delta-i\frac{\Gamma}{2}\right)P_{e}-\frac{\Omega}{2}\left(\bm{\mathcal{\epsilon}}(\hat{x})\cdot\bm{\hat{D}}^{\dagger}+h.c.\right).$ (3)
@@
The atomic operator $\hat{\bm{D}}^{\dagger}$ is a raising operator for all the hyperfine levels in the ground-state, and $\hat{\bm{D}}$ is the corresponding lowering operator. They are defined in terms of the dipole operator $\bm{\hat{d}}$ normalized by the reduced matrix element,
@@
The Hamiltonian in Eq. 3 also contains a non-Hermitian imaginary decay rate $\Gamma$ of the excited state due to spontaneous emission. In the Wigner-Weisskopf model, the excited state decays due to spontaneous emission without repopulating the ground-state. This leads to a reduction in the wavefunction normalization. The master equation for the density matrix contains refeeding terms with collapse operators $L_{i}$ that repopulate the ground-state and preserve the density matrix normalization:
$\frac{d\hat{\rho}}{dt}=-i(\hat{H}\hat{\rho}-\hat{\rho}\hat{H}^{\dagger})+\sum_{i}\hat{L}_{i}\hat{\rho}\hat{L}^{\dagger}_{i}.$ (6)
@@
$\hat{L}_{\hat{\bm{n}},q}=\sqrt{\Gamma}\,\hat{R}_{\bm{n},q}\hat{D}_{q}.$ (7)
The collapse operators have two effects when operating on the density matrix. First, they lower the atom back to the ground-state. $\hat{D}_{q}$ is the lowering operator in the spherical vector basis $\hat{D}_{q}=\hat{\bm{e}}^{*}_{q}\cdot\hat{\bm{D}}$, where $\hat{\bm{e}}_{q}$ is a spherical basis vector. Second, the recoil operator $\hat{R}_{n,q}$ imparts a momentum kick due to the photon recoil after the emission of a photon with polarization $q$ in $\bm{n}$ directions. In general, the recoil operator is averaged over the angular emission profile for each polarization. If we assume that the photon is emitted along either direction of the $x$-axis, then the recoil operators are $R_{\pm,q}=e^{\pm ikx}$, and we have a total of six collapse operators for the two emissions directions and three polarizations
$\hat{L}_{\pm,q}=\sqrt{\Gamma}\,e^{\pm ik\hat{x}}\hat{D}_{q}.$ (8)
In 1D, the photon recoil operator $R_{\pm}$ = $e^{\pm ik\hat{x}}$ can be understood as a momentum translation operator that induces a momentum kick due to the recoil from an emitted photon, with $e^{\pm ik\hat{x}}|\psi(p)\rangle=|\psi(p\pm\hbar k)\rangle$.
We can express the position operator in terms of the harmonic oscillator creation and annihilation operators $\hat{x}=x_{0}(\hat{a}+\hat{a}^{\dagger})$, where $x_{0}=\sqrt{\hbar/2m\nu}$ is the width of the ground-state wavefunction. The recoil operator becomes $e^{ik\hat{x}}=e^{i\eta(\hat{a}+\hat{a}^{\dagger})}$, where the Lamb-Dicke (LD) parameter $\eta=kx_{0}$ is the ratio of the ground-state wavefunction width to the wavelength.
The square of the LD parameter also happens to be equal to the ratio of the photon recoil energy $E_{\text{recoil}}=\hbar^{2}k^{2}/(2m)$ to the harmonic energy $\hbar\nu$, or $\eta^{2}=E_{\text{recoil}}/(\hbar\nu)$. When $\eta\ll 1$, then the recoil from photon emission or absorption is unlikely to change $n$, and the ground-state wavefunction is smaller than the cooling laser’s wavelength. In this regime, we can also approximate the recoil operator as
@@
The result of a recoil is then $e^{ik\hat{x}}|n\rangle\approx|n\rangle+\eta\sqrt{n}|n-1\rangle+\eta\sqrt{n+1}|n+1\rangle$. The zeroth order term is from $n\rightarrow n$, and the first order LD term is from $n\rightarrow n\pm 1$. The LD regime is defined as when $\eta^{2}(2n+1)\ll 1$. One common misconception is that in the LD regime, heating is suppressed. Although the probability of $n$ changing decreases, the overall heating rate is still given by the scattering rate times the recoil energy.
More generally, the matrix element can be expressed in terms of the generalized Laguerre polynomial as [36]
@@
Here, $n_{<}$ is the least of $n$ and $n^{\prime}$, and $\Delta n=|n^{\prime}-n|.$ This functional form becomes important when cooling outside the LD regime. The $n\rightarrow n-1$ transition goes to zero for higher $n$, creating dark motional states where the population gets stuck. Cooling can still occur by addressing higher order sidebands, as shown experimentally in Ref. [26].
The master equation here is simulated throughout this paper for different cooling schemes using the steady-state solver in QuTip [37]. We include up to 15 harmonic states, which for the larger spin systems in this paper like $F=1$ to $F=2$ takes approximately one minute per simulation.
II.2 Effective master equation within ground-state subspace
Most optical cooling processes take place when there is low saturation, meaning that only a small amount of the steady-state population is in the excited state. The excited state can then be reasonably approximated from the ground state, and can be adiabatically eliminated. In this section, we will derive the effective ground-state master equation.
In situations where the excited state quickly reaches an equilibrium, the excited state can be solved in terms of the ground-states and then eliminated from the time evolution equations. The elimination results in an effective Hamiltonian and collapse operator that gives the dynamics within the ground-state. These effective operators can be derived by performing these operations for a given Hamiltonian [38], but Ref. [39] provides a compact general formalism that we use here.
This first form is valid even when the cooling light is close to resonance such as in sideband cooling. The Hamiltonian and collapse operators within the ground-state subspace are
@@ II.2 Effective master equation within ground-state subspace
Here the laser detuning is $\Delta_{m_{g},m_{e}}=\omega_{L}-(E_{m_{e}}-E_{m_{g}})$ appears in the denominator. For the case of an atom with spin interacting with a single frequency laser field, as in the Hamiltonian in Eq. 3, the effective ground-state Hamiltonian is
@@
[…]
and the collapse operator for the emission of a photon with polarization $q$ is:
$L_{\text{eff}}^{q,\bm{n}}=\!-\sqrt{\Gamma}\,\frac{\Omega}{2}\!\!\sum_{m_{g},m_{e}}\!\!\frac{(\mathcal{R}_{\bm{n}}\bm{e}_{q}^{*}\cdot\bm{D})|m_{e}\rangle\langle m_{e}|(\bm{\epsilon}(\hat{x})\cdot\bm{D}^{\dagger})|m_{g}\rangle\langle m_{g}|}{\Delta_{m_{g},m_{e}}-i\Gamma/2}.$ (14)
The effective Hamiltonian includes light shifts, two-photon Raman coupling, and decay due to spontaneous emission. The collapse operators repopulate the ground-state.
Two photon cooling schemes operate in the large detuning limit. In the large detuning limit, the state dependence in the denominator detuning vanishes $\Delta_{m_{g},m_{e}}\approx\Delta$. The sum over the excited states in the numerator becomes the identity. The effective ground-state operators simplify to:
@@
For the single frequency Hamiltonian in Eq. 3, it becomes
@@
and the ground-state collapse operator in the large detuning limit is
$L_{\text{eff}}^{q,\bm{n}}=-\sqrt{\Gamma}\,\frac{\Omega}{2}\left[\frac{(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*}\cdot\bm{D})(\bm{\epsilon}(\hat{x})\cdot\bm{D}^{\dagger})}{\Delta-i\Gamma/2}\right].$ (18)
The interaction part of the ground-state Hamiltonian describes two-photon processes such as light shifts and Raman transitions. The right term in the numerator $\bm{\epsilon}(\hat{x})\cdot\hat{\bm{D}}^{\dagger}$ excites the atom due to absorption of a photon from the cooling light. The left term $\bm{\epsilon}^{*}(\hat{x})\cdot\hat{\bm{D}}$ returns the atom to the ground-state with the emission of a photon into the cooling light.
The ground-state collapse operator describes the redistribution of population within the ground-state due to spontaneous emission. It has a similar form compared to interaction Hamiltonian. The right term $\bm{\epsilon}(\hat{x})\cdot\hat{\bm{D}}^{\dagger}$ excites the atom due to the absorption of a photon from the cooling light. The left term $\mathcal{R}_{\bm{n}}\bm{e}_{q}^{*}\cdot\hat{\bm{D}}$ returns the atom to the ground-state by emitting a photon into free-space and gives the atom a momentum kick through the recoil operator $\mathcal{R}_{\bm{n}}$. This momentum kick is the key difference between the two operators.
After adiabatic elimination, the form of both the effective Hamiltonian and collapse operators are in terms of the scalar products of the dipole operators and field polarization,
$H\propto(\bm{A}^{*}\cdot\hat{\bm{D}})(\hat{\bm{D}}^{\dagger}\cdot\bm{B})=(\bm{A}^{*}\otimes\bm{B})\cdot(\hat{\bm{D}}\otimes\hat{\bm{D}}^{\dagger}).$ (19)
These operators act within the ground-state spin manifold just like the hyperfine spin operators $\hat{F}_{i}$, $F_{\pm}$. If you observe, the right hand side is actually a scalar product of two rank 2 matrices. Using the Wigner-Eckart theorem and properties of spherical vectors, we can express these in terms of spin operators. The spin Hamiltonian is [38]
@@
$\left.+C^{(1)}i(\bm{\epsilon}^{*}(\hat{x})\times\bm{\epsilon}(\hat{x}))\right.$
$\left.+C^{(2)}\left(((\bm{\epsilon}^{*}(\hat{x})\!\cdot\!\hat{\bm{F}})(\bm{\epsilon}(\hat{x})\!\cdot\!\hat{\bm{F}})+c.c.)/2-\frac{1}{3}\hat{\bm{F}}^{2}|\bm{\epsilon}(\hat{x})|^{2}\right)\right]$
and the collapse operators which give the redistribution of population in the ground-state due to spontaneous emission are
$L_{\text{eff}}=-\sqrt{\Gamma}\frac{\Omega}{2\Delta}\left[C^{(0)}(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\cdot\bm{\epsilon}(\hat{x})\right.$ (21)
$\left.+C^{(1)}i((\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\times\bm{\epsilon}(\hat{x}))\right.$
$\left.+\frac{C^{(2)}}{2}\left[((\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\!\cdot\!\hat{\bm{F}})(\bm{\epsilon}(\hat{x})\!\cdot\!\hat{\bm{F}})+c.c.\right)\right.$

### v3

@@ Abstract
Cooling atoms to the ground state of optical tweezers is becoming increasingly important for high-fidelity imaging, cooling, and molecular assembly. While extensive theoretical work has been conducted on cooling in free space, fewer studies have focused on cooling in bound states. In this work, we present a unified formalism for optical cooling mechanisms in neutral atom tweezers, including resolved and unresolved sideband cooling with different trapping potentials, polarization gradient cooling, gray molasses cooling, $\Lambda$-enhanced gray molasses cooling, and Raman sideband cooling. We perform simulations with hyperfine levels and demonstrate good agreement with a simplified spin model. We derive and discuss the fundamental limits of each cooling mechanism and propose new strategies for achieving ground state cooling in optical tweezers. Our findings provide valuable insights into optimizing cooling schemes for neutral atoms in optical tweezers, paving the way for minimizing thermal decoherence in Rydberg and molecular gates and improving efficiencies of molecular assembly.
@@ I Introduction
Laser cooling of bound neutral atoms has recently gained renewed significance in experiments involving tweezer arrays of atoms and molecules [1]. In systems of Rydberg atoms [2] or molecules [3, 4, 5] with dipolar coupling, thermal motion limits the coherence of interactions. Molecular assembly [6, 7] and few-body physics [8, 9] studies require high-fidelity ground state preparation of multiple atoms.
@@
A wide range of optical cooling techniques have been developed for trapped atoms, illustrated in Fig. 1. Single-photon cooling schemes, such as sideband cooling, have been used for atoms with narrow lines for high-fidelity imaging and ground state preparation [10, 11, 12]. Two-photon schemes, including polarization gradient (PG) [13, 14, 15, 16], gray molasses (GM) and $\Lambda$-enhanced gray molasses ($\Lambda$-GM) [17, 18, 19, 20, 21], and Raman-sideband cooling (RSB) [22, 23, 24, 25, 26], have been successfully applied for imaging and cooling of atoms without narrow lines. While extensive theoretical work has been developed for sideband cooling [27, 28, 29], many of the other cooling schemes for bound atoms are still interpreted through their free-space pictures, for example in Sisyphus cooling where an atom moves through a changing polarization [13]. However, this picture breaks down for a tightly trapped atom in the Lamb-Dicke regime, where the atom wavefunction is smaller than the wavelength.
@@
In this work, we develop a generalized optical cooling model for a bound atom with spin in a single or counter-propagating beam configuration with different polarizations. We validate these models using master equation simulations that include the ground and excited spins and the harmonic oscillator states. Our analysis reveals a consistent picture across various cooling schemes, shown in Fig. 1. Cooling occurs when $|n\rangle$ and $|n-1\rangle$ are Raman coupled for different spin states and brought into resonance using a combination of light shifts (PG, GM, $\Lambda$-GM) and two-photon detunings (RSB).
@@
Sideband cooling has been extensively studied in previous works [30, 31], but studies for cooling a bound atom with spin has been much more limited. Previous theoretical work on cooling bound atoms with spin has primarily focused on specific cases, such as $J\!\!=\!\!1/2$ to $J^{\prime}\!\!=\!\!3/2$ transitions in counter-propagating orthogonal linear light [32] or in a linear light standing wave [33]. We develop a new formalism for arbitrary spin and polarization, with a particular emphasis on the experimentally relevant $\sigma_{+}\!-\!\sigma_{-}$ configuration. Notably, while these previous theoretical studies found a fundamental limit of $\langle n\rangle\approx 1$, we demonstrate that schemes like GM and $\Lambda$-GM can be used for high ground state preparation when considering different polarization and spin configurations.
@@
Figure 1: (a) Laser cooling of a trapped atom with ground and excited state spin $F$ and $F^{\prime}$ in a changing polarization from counter-propagating beams. (b) Single photon cooling schemes such as resolved and unresolved sideband cooling drive population from the ground state to the excited state with one reduced motional quanta $n-1$. In the Lamb-Dicke regime, the recoil heating is smaller than the trapping frequency and the population mostly returns to $n-1$, resulting in net cooling. (c) Two-photon cooling schemes like polarization gradient (PG), gray molasses (GM), and $\Lambda$-enhanced gray molasses ($\Lambda$-GM) drive $n\rightarrow n-1$ between two ground state spins with a two-photon Raman transition. (d) Temperature scale for cooling schemes. A wide variety of cooling schemes has been used to trap, cool, and image tightly trapped atoms.
@@
A recent experimental result achieved a record-high imaging fidelity of 99.96% for lithium using $\Lambda$-enhanced GM cooling with significant ground state preparation, despite lithium having the largest recoil heating [21]. This raises the question of whether spin cooling techniques can be further optimized for ground state preparation. The unified model also allows us to explore novel cooling schemes that transcend the conventions of any single conventional cooling technique.
@@
The paper is structured as follows. Section II develops the general master equation for spin cooling and derives effective ground state spin operators by adiabatically eliminating the excited state. Section III applies this formalism to sideband cooling, including cases with different excited state trapping frequencies. Section IV develops a formalism for spin cooling and applies it to PG, GM, $\Lambda$-GM, and RSB cooling. Each section compares exact simulations to simplified models and contrasts various cooling techniques, with the goal of creating a unified picture that will lead to more coherent interactions and higher fidelity ground state preparation.
@@ II.1 Master equation for laser cooling
In this section, we develop a master equation for an atom confined in a tight harmonic trap and interacting with a semi-classical electric field. In a semi-classical approximation, we take the expectation of the electric field but keep the atom position $\hat{x}$ as a quantum operator [34, 35]. The electric field for a single laser frequency $\omega_{L}$ is:
$\bm{E}(\hat{x},t)=E_{0}\,{\bm{\mathcal{\epsilon}}}(\hat{x})\,e^{-i\omega_{L}t}+h.c.$ (1)
$E_{0}$ is a real field amplitude and the last term is the hermitian conjugate. The complex polarization and phase are contained within $\bm{\epsilon}(\hat{x})$ and can describe a traveling wave or a standing wave with arbitrary polarization configurations.
@@
$H=\nu\hat{a}^{\dagger}\hat{a}+\left(\Delta-i\frac{\Gamma}{2}\right)P_{e}-\frac{\Omega}{2}\left(\bm{\mathcal{\epsilon}}(\hat{x})\cdot\bm{\hat{D}}^{\dagger}+h.c.\right).$ (3)
@@
The atomic operator $\hat{\bm{D}}^{\dagger}$ is a raising operator for all the hyperfine levels in the ground state to the excited state, and $\hat{\bm{D}}$ is the corresponding lowering operator. They are defined in terms of the dipole operator $\bm{\hat{d}}$ normalized by the reduced matrix element,
@@
The Hamiltonian in Eq. 3 also contains a non-Hermitian imaginary decay rate $\Gamma$ of the excited state due to spontaneous emission. In the Wigner-Weisskopf model, the excited state decays due to spontaneous emission without repopulating the ground state. This leads to a reduction in the wavefunction normalization. The master equation for the density matrix contains refeeding terms with collapse operators $L_{i}$ that repopulate the ground state and preserve the density matrix normalization:
$\frac{d\hat{\rho}}{dt}=-i(H\hat{\rho}-\hat{\rho}H^{\dagger})+\sum_{i}{L}_{i}\hat{\rho}{L}^{\dagger}_{i}.$ (6)
@@
${L}_{\hat{\bm{n}},q}=\sqrt{\Gamma}\,\hat{R}_{\hat{\bm{n}},q}\hat{D}_{q}.$ (7)
The collapse operators have two effects when operating on the density matrix. First, they lower the atom back to the ground state. $\hat{D}_{q}$ is the lowering operator in the spherical vector basis $\hat{D}_{q}=\hat{\bm{e}}^{*}_{q}\cdot\hat{\bm{D}}$, where $\hat{\bm{e}}_{q}$ is a spherical basis vector. Second, the recoil operator $\hat{R}_{n,q}$ imparts a momentum kick due to the photon recoil after the emission of a photon with polarization $q$ in direction $\hat{\bm{n}}$. In general, the recoil operator is averaged over the angular emission profile for each polarization. If we assume that the photon is emitted along either direction of the $x$-axis, then the recoil operators are $R_{\pm,q}=e^{\pm ikx}$, and we have a total of six collapse operators for the two emissions directions and three polarizations
${L}_{\pm,q}=\sqrt{\Gamma}\,e^{\pm ik\hat{x}}\hat{D}_{q}.$ (8)
In 1D, the photon recoil operator can be understood as a momentum translation operator that induces a momentum kick due to the recoil from an emitted photon, with $e^{\pm ik\hat{x}}|\psi(p)\rangle=|\psi(p\pm\hbar k)\rangle$.
We can express the position operator in terms of the harmonic oscillator creation and annihilation operators $\hat{x}=x_{0}(\hat{a}+\hat{a}^{\dagger})$, where $x_{0}=\sqrt{\hbar/2m\nu}$ is the width of the ground state wavefunction. The recoil operator becomes $e^{ik\hat{x}}=e^{i\eta(\hat{a}+\hat{a}^{\dagger})}$, where the Lamb-Dicke (LD) parameter $\eta=kx_{0}$ is the ratio of the ground state wavefunction width to the wavelength of the cooling light.
The square of the LD parameter also happens to be equal to the ratio of the photon recoil energy $E_{\text{recoil}}=\hbar^{2}k^{2}/(2m)$ to the harmonic energy $\hbar\nu$, or $\eta^{2}=E_{\text{recoil}}/(\hbar\nu)$. When $\eta\ll 1$, then the recoil from photon emission or absorption is unlikely to change $n$, in addition to the ground state function being smaller than the cooling laser’s wavelength. One common misconception is that in the LD regime, heating is suppressed. Although the probability of $n$ changing decreases, the overall heating rate is still given by the scattering rate times the recoil energy. In this regime, we can also approximate the recoil operator to first order in the LD parameter as
@@
The result of a recoil is then to transfer population to neighboring motional states
$e^{ik\hat{x}}|n\rangle\approx|n\rangle+\eta\sqrt{n}|n-1\rangle+\eta\sqrt{n+1}|n+1\rangle$ (10)
The zeroth order term is from $n\rightarrow n$, and the first order LD term is from $n\rightarrow n\pm 1$. The LD regime is defined as when $\eta^{2}(2n+1)\ll 1$.
Outside the LD regime, the matrix element of the recoil operator between different motional states can be expressed in terms of the generalized Laguerre polynomial as [36]
@@
Here, $n_{<}$ is the least of $n$ and $n^{\prime}$, and $\Delta n=|n^{\prime}-n|.$ This functional form becomes important when cooling outside the LD regime. The $n\rightarrow n-1$ transition goes to zero for higher $n$, creating dark motional states where the population gets stuck. Cooling can still occur by addressing higher order sidebands, as shown in Ref. [37, 38, 26].
The master equation here is simulated throughout this paper for different cooling schemes using the steady-state solver in QuTip [39]. We include up to 15 harmonic states, which for the larger spin systems in this paper like $F=1$ to $F^{\prime}=2$ takes approximately one minute per simulation.
II.2 Effective master equation within ground state subspace
Optical cooling typically takes place in the low saturation regime, meaning that only a small amount of the steady-state population is in the excited state. The excited state can be accurately approximated from the ground state, and can be adiabatically eliminated to create effective ground state dynamics. In this section, we will derive the effective ground state master equation.
In situations where the excited state quickly reaches an equilibrium, the excited state can be solved in terms of the ground states and then eliminated from the time evolution equations. The elimination results in an effective Hamiltonian and collapse operator that gives the dynamics within the ground state. These effective operators can be derived by performing these operations for a given Hamiltonian [40], but Ref. [41] provides a compact general formalism that we use here.
Most generally, for light that is either close to resonance or far-from resonance, the effective ground state Hamiltonian and collapse operators can be expressed as a sum over the ground and excited spin states as [41]
@@ II.2 Effective master equation within ground state subspace
The laser detuning between two specific spin states is $\Delta_{m_{g},m_{e}}=\omega_{L}-(E_{m_{e}}-E_{m_{g}})$. The first and last operators in both of these expressions operate only on ground state spins.
Substituting the Hamiltonian for a spin interacting with a single laser frequency from Eq. 3 gives an effective Hamiltonian
@@
$\frac{\Omega^{2}}{8}\!\left[\sum_{m_{g},m_{e}}\!\!\frac{(\bm{\epsilon}(\hat{x})\!\cdot\!\bm{D})|m_{e}\rangle\langle m_{e}|(\bm{\epsilon}^{*}(\hat{x})\!\cdot\!\bm{D}^{\dagger})|m_{g}\rangle\langle m_{g}|}{\Delta_{m_{g},m_{e}}-i\Gamma/2}+h.c.\right]$
and the collapse operator for the emission of a photon with polarization $q$ becomes
$L_{\text{eff}}^{q,\hat{\bm{n}}}=\!-\sqrt{\Gamma}\,\frac{\Omega}{2}\!\!\sum_{m_{g},m_{e}}\!\!\frac{(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*}\cdot\bm{D})|m_{e}\rangle\langle m_{e}|(\bm{\epsilon}(\hat{x})\cdot\bm{D}^{\dagger})|m_{g}\rangle\langle m_{g}|}{\Delta_{m_{g},m_{e}}-i\Gamma/2}.$ (15)
The effective Hamiltonian includes light shifts, two-photon Raman coupling, and decay due to spontaneous emission. The collapse operators repopulate the ground state after spontaneous emission. These expressions are valid close to resonance and will be used for resolved and unresolved sideband cooling.
Two photon cooling schemes typically operate in the large detuning limit. In the large detuning limit, the state-dependent detuning can be replaced by a constant detuning $\Delta_{m_{g},m_{e}}\approx\Delta$. The sum over the excited states in the numerator in both operators becomes the identity. The effective ground state operators from Eqs. 12 and 13 simplify to
@@
The single frequency spin operators from Eqs. 14 and 15 become
@@
$L_{\text{eff}}^{q,\hat{\bm{n}}}=-\sqrt{\Gamma}\,\frac{\Omega}{2}\left[\frac{(\mathcal{R}_{\hat{\bm{n}}}\hat{\bm{e}}_{q}^{*}\cdot\bm{D})(\bm{\epsilon}(\hat{x})\cdot\bm{D}^{\dagger})}{\Delta-i\Gamma/2}\right].$ (19)
The ground state Hamiltonian describes two-photon processes such as light shifts and Raman transitions. The right term in the numerator $\bm{\epsilon}(\hat{x})\cdot\hat{\bm{D}}^{\dagger}$ excites the atom due to absorption of a photon from the cooling light. The left term $\bm{\epsilon}^{*}(\hat{x})\cdot\hat{\bm{D}}$ returns the atom to the ground state with the emission of a photon into the cooling light.
The ground state collapse operator describes the redistribution of population within the ground state due to spontaneous emission. It has a similar form as the interaction Hamiltonian. The right term $\bm{\epsilon}(\hat{x})\cdot\hat{\bm{D}}^{\dagger}$ excites the atom due to the absorption of a photon from the cooling light. The left term $\mathcal{R}_{\bm{n}}\bm{e}_{q}^{*}\cdot\hat{\bm{D}}$ returns the atom to the ground state by emitting a photon into free-space and gives the atom a momentum kick through the recoil operator $\mathcal{R}_{\bm{n}}$.
The operators operate within the ground state spin manifold. Therefore, we can also express these operators in terms of the spin operators $\hat{F}_{i}$. While the derivation is long, the idea is quite simple. After adiabatic elimination, both operators are in the form
$(\bm{A}^{*}\cdot\hat{\bm{D}})(\bm{B}\cdot\hat{\bm{D}}^{\dagger})=(\bm{A}^{*}\otimes\bm{B})\cdot(\hat{\bm{D}}\otimes\hat{\bm{D}}^{\dagger}).$ (20)
Here $\bm{A}$ and $\bm{B}$ are vectors related to the electric field of the absorbed or emitted photons. On the right side we re-express it as a scalar product of two rank 2 tensors. The operator $\hat{\bm{D}}\otimes\hat{\bm{D}}^{\dagger}$ is a rank 2 tensor, formed by the two tensor product of two rank 1 vectors. This rank 2 operator can be decomposed into an irreducible representation which consists of a rank 0 scalar, rank 1 vector, and rank 2 symmetric and traceless tensor. Finally, the Wigner-Eckert theorem can be used on each of these components to convert them to spin matrix elements. The spin Hamiltonian becomes [40]
@@
$\left.+C^{(1)}i[\bm{\epsilon}^{*}(\hat{x})\times\bm{\epsilon}(\hat{x})]\cdot\hat{\bm{F}}\right.$
$\left.+C^{(2)}\left(\frac{1}{2}[(\bm{\epsilon}(\hat{x})\!\cdot\!\hat{\bm{F}})^{\dagger}(\bm{\epsilon}(\hat{x})\!\cdot\!\hat{\bm{F}})+c.c.]-\frac{1}{3}\hat{\bm{F}}^{2}|\bm{\epsilon}(\hat{x})|^{2}\right)\right].$
The coefficients $C^{(i)}$ are given in an appendix. The collapse operators which give the redistribution of population in the ground state due to spontaneous emission are
$L_{\text{eff}}=-\sqrt{\Gamma}\frac{\Omega}{2(\Delta-i\Gamma/2)}\left[C^{(0)}(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\cdot\bm{\epsilon}(\hat{x})\right.$ (22)
$\left.+C^{(1)}i([(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\times\bm{\epsilon}(\hat{x})]\cdot\hat{\bm{F}})\right.$
$+C^{(2)}\left[\frac{1}{2}[(\mathcal{R}_{\bm{n}}\hat{\bm{e}}_{q}^{*})\!\cdot\!\hat{\bm{F}})^{\dagger}(\bm{\epsilon}(\hat{x})\!\cdot\!\hat{\bm{F}})+h.c.]\right.$